OpenStudy (nuccioreggie):

The equation shows the relationship between x and y: y = 7x + 2 What is the slope of the equation? −7 −5 2 7

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (trojanpoem):

y = mx + b ( b: y-intersection, m: slope) y = 7x + 2 Guess the slope.

OpenStudy (nuccioreggie):

i think its C

OpenStudy (texaschic101):

the slope will be the number before the x

OpenStudy (trojanpoem):

m = 7 , b= 2

OpenStudy (nuccioreggie):

so its B?

OpenStudy (trojanpoem):

Nah, D

OpenStudy (trojanpoem):

y = -x + 5 (-1 is the slope) 5 y = 6x + 10 y = 6/5 x + 2 (6/5 is the slope) and so on

OpenStudy (nuccioreggie):

oh ok ty can you help me with like a couple more please